OpenStudy (anonymous):

A rectangle is 5 feet longer than it is wide. The perimeter of the rectangle is 34 feet. What is the length of the rectangle?

OpenStudy (anonymous):

\[P = 2(L+W)\] \[34 = 2(2W + 5)\] \[12 = 2W\] \[W = 6\] \[L = 11\]

OpenStudy (anonymous):

2L+2W=34 L=W+5 2L+2(L-5)=34 4L=44 L=11